Grade 12Discuss with colleagues and IITians

two circles x2+y2-6x+8=0 and x2+y2-6=0 are given.The eqn. of circle through their point of intersection and (1,1) is?

please explain!!!!

i will surely approve

pratiksha agarwal
13 Years agoGrade 12

5 Answers

Ashwin Muralidharan IIT Madras
ApprovedApproved Tutor Answer13 Years ago

On solving,

You should be getting k = -1/2

and eqaution of circle as x^2 + y^2 - 3x + 1 = 0.

 

Regards,

Ashwin (IIT Madras).

FITJEE
ApprovedApproved Tutor Answer13 Years ago
step1----
             get the two intersection                ofthe two circles by                      solving the two                            equations.                          
step 2---- now U got three pts                     from where that circle is               passing i.e. two                           intersectn points+(1,1).
step3----
             use the formula----
       
      (x-x1)(x-x2)+(y-y1)(y-y2)=0
     and pass it from third one.
     where,
              (x1,y1)
              (x2,x3)
              (x3,y3)
        are the three points from             where the circle is passing.
   

Vikas TU
ApprovedApproved Tutor Answer13 Years ago

eqn. of circle through their point of intersectionis the case of the family of circles.

and the condiotion is =

S1 + KS2 = 0

WHERE S1 and S2  =  eqn. of circles

and k is any constant.

Thus,

(x2+y2-6x+8) + k(x2+y2-6) = 0......................................(1)

Now pass the points (1,1) in the eqn. (1)

U will get the value of K.

After then put the value of k  again and get the eqn. of circle.

 


somy chaturvedi
ApprovedApproved Tutor Answer13 Years ago

x^2+y^2-3x+1=0.if i am right then i willl send u the solution

FITJEE
13 Years ago

step1----
             get the two intersection                ofthe two circles by                      solving the two                            equations.                          
step 2---- now U got three pts                     from where that circle is               passing i.e. two                           intersectn points+(1,1).
step3----
             use the formula----
       
      (x-x1)(x-x2)+(y-y1)(y-y2)=0
     and pass it from third one.
     where,
              (x1,y1)
              (x2,x3)
              (x3,y3)
        are the three points from             where the circle is passing.